OpenStudy (cheesecakekitten):

WILL GIVE MEDAL TO BEST AND SHORTEST ANSWER The system of equations is coincident. What are the missing values? x +3y= 5 4x + (blank)y = (blank)

OpenStudy (cheesecakekitten):

PLZ ANSWER QUICK

OpenStudy (radar):

Do like they started, multiply the first equation thru by 4-------and make it quick!

OpenStudy (cheesecakekitten):

i guess i will try :P

OpenStudy (cheesecakekitten):

google calculator won't do variable equations -.-

OpenStudy (cheesecakekitten):

...

OpenStudy (radar):

OK they multiplied the x by 4 you continue and multiply the 3y by 4 (you can do this in your head) Now continue on and multiply the 5 by 4, again no calculator needed.

OpenStudy (radar):

Times up!

OpenStudy (radar):

Do this: 4(x+3y) = 4 (5) Good luck with your studies, bye.

OpenStudy (anonymous):

4(x+3y) = 4 (5) is correct because i am in algebra 3

OpenStudy (cheesecakekitten):

Okay, if you're sure :P

OpenStudy (cheesecakekitten):

Bye, thanks

OpenStudy (anonymous):

welcome